OpenStudy (anonymous):

Use the table below to evaluate: d/dx[f[g(2x)]] at x=1. x 1 2 3 4 f(x) 6 1 8 2 f ′(x) 6 1 8 2 g(x) 1 4 4 3 g ′(x) 9 5 5 –4

OpenStudy (anonymous):

@welshfella @Fzxt

OpenStudy (adamaero):

Think of it as a nest (since it is nested functions) Start in the middle of the nest, and work outward... Show what you get. I'll start: 2*x = 2*1 g(2) = 4 then 4 into...next layer...

OpenStudy (anonymous):

I'm not quite sure I understand what you're saying